# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ E.3.1 Layers II
Figure 7: Total energy transfer from the 3D waves to the 2D condensate as a function of $S\equiv{U^{\prime}}/\Omega$. Integration is carried out without homochiral-wave filter $F_{k_{f}}$ (light blue), and with a Lorentzian form (140) (blue), both without cutoff, $K\to\infty$. Dashed line: leading order behavior (104). Red line: mean-flow dissipation $\nu{U^{\prime}}^{2}=\frac{S^{2}}{4Ro_{\epsilon}^{2}}$ for various values of $Ro_{\epsilon}$. The solution to the mean-wave system is found at the intersection between the blue and the red line (black square). The flux-loop state $T_{\rm 3D-2D}=0$, obtained when $Ro_{\epsilon}\to\infty$, is marked as a green square.
@@ E.4 Numerical integration without cutoff
$F_{k_{f}}=\frac{1}{\left(1+\left(\frac{1}{S}Q\frac{4\pi p_{z}\sqrt{k_{f}^{2}-p_{x}^{2}-p_{z}^{2}}}{k_{f}^{3}}\right)^{2}\right)^{1/2}},$ (140)
@@
Note, finally, that we have neglected 2D-2D interactions in this computation. With isotropic forcing, these should lead to an energy input $T_{\rm 2D-2D}\simeq l_{f}/(4L_{z})\ll 1$ (see [gome2025waves]), which is negligible for sufficiently large values of $Ro_{\epsilon}$, but not too large. With this hypothesis, the energy injection rate $\epsilon$ corresponds to the energy input in 3D modes only. At low rotation, considering 2D-2D interactions goes beyond our QL theory, as they can disappear due to instabilities, which are not included in our framework.
@@ F.1 Choice of cutoff
The eddy turnover time $\tau_{nl}^{\rm eddy}$ is determined from the Kolmgorov scaling, while the wave-wave interaction time scale $\tau_{nl}^{\rm wave}$ is predicted from wave-turbulent theory [galtier2003weak].
@@
$k_{U}^{\rm wave}$ $=(\Omega p_{z})^{1/3}S^{2/3}\epsilon^{-1/3},~~~\text{the crossover from condensate- to wave-dominated interactions}$ (157)
$k_{U}^{\rm eddy}$ $=(\Omega S)^{3/2}\epsilon^{-3/2},~~~\text{the crossover from condensate- to eddy-dominated interactions,}$ (158)
with $S\equiv\frac{{U^{\prime}}}{\Omega}$. These various wavenumbers are explained in the following.
@@
Nonlinear interactions between fluctuation modes occur over a nonlinear time scale $\tau_{nl}$, whose estimation depends on whether mode $\bm{p}$ is wave-dominated ($p<k_{\Omega}$) or eddy-dominated ($p>k_{\Omega}$). When modes are wave-dominated, they interact with other waves over a slow nonlinear timescale $\tau_{nl}^{\rm wave}$ in (155), which is $\gg\tau_{\Omega}$. In contrast, when modes are eddy-dominated, they interact with each other over the Kolmogorov time scale $\tau_{nl}^{\rm eddy}$ in (154).
@@ F.2 Numerical integration of the Reynolds stress with cutoff
When integrating numerically the Reynolds stress, we consider the Lorentzian filter for the homochiral-wave resonances (140), with parameter $Q=8$. It yields a smooth transition between the homochiral regime and the mixed-interaction regime around $Ro_{\epsilon}\simeq 0.1$ (see Fig. 9). This near-resonant treatment is more accurate at intermediate $Ro_{\epsilon}$, as it captures the contribution of small near-resonances. However, it fails in the limit $Ro_{\epsilon}\to 0$, as explained in [gome2025waves].
@@
$\omega_{\bm{p}}^{+}+\omega_{\bm{q}}^{-}<\tilde{\omega}{U^{\prime}}$ (162)
in place of (37). This is equivalent to changing $S$ to $\tilde{\omega}S$ in (113). Parameter $\tilde{\omega}$, which must be $\sim 1$, provides an adaptable broadening of the resonant condition. It corresponds to approximating the oscillating factor $\Delta(t)^{s,-s}$ (33) with a parametrized Heaviside function $\Theta(1-|\omega_{\bm{p}\bm{q}}^{+-}|/(\tilde{\omega}{U^{\prime}})$. In the results established so far, and in the solution shown in Fig. 4 of the main text, $\tilde{\omega}$ is set to 1. In general, $\tilde{\omega}$ can be used as a tunable parameter of the numerical integration which can better fit the DNS data points. In Figs. 9(a,b), we use $\tilde{\omega}=1.2$. Solutions with $\tilde{\omega}=1$ and $\tilde{\omega}=1.2$ are compared in Fig. 9(c) and (d), respectively.
@@
Figure 10: Energy transfer to the 2D flow, $T_{\rm 3D-2D}/\epsilon$, as a function of RoRo (a) and $Ro_{\epsilon}$ (b). The dashed line in (b) shows the analytical solution without cutoff, Eq. (152)
The QL computation involving cutoff $K$ is compared to the DNS data points in Figs. 9(c-d). The agreement is quantitatively excellent up to $Ro_{\epsilon}\simeq 0.5$. At large $Ro_{\epsilon}$, the behavior is sensitive to the cutoff and ${U^{\prime}}/\Omega$ decays due to a wave shortage, and reaches zero at $Ro\simeq 0.5$. Our QL result qualitatively reproduces the decay of ${U^{\prime}}/\Omega$ at fixed $Re$, and the transition to strong turbulence seen in the DNS. In addition, we obtain a close quantitative agreement at $Re=23$ and $46$ when using parameters $\tilde{\omega}=1.2$ and $Q=8$. (Note that we obtain also a good quantitative agreement with the non-cutoff case $\tilde{\omega}=1$ below $Ro_{\epsilon}=1$).

### v2

@@ E.3.1 Layers II
Figure 7: Total energy transfer from the 3D waves to the 2D condensate as a function of $S\equiv{U^{\prime}}/\Omega$. Integration is carried out without homochiral-wave filter $F_{k_{f}}$ (light blue), and with a Lorentzian filter (140) (blue), both without cutoff ($K\to\infty$). Dashed line: leading order behavior (104). Red line: mean-flow dissipation $\nu{U^{\prime}}^{2}=\frac{S^{2}}{4Ro_{\epsilon}^{2}}$ for various values of $Ro_{\epsilon}$. The solution to the mean-wave system is found at the intersection between the blue and the red line (black squares). The flux-loop state $T_{\rm 3D-2D}=0$, obtained when $Ro_{\epsilon}\to\infty$, is marked as a green square. The yellow region $S<l_{f}/L_{z}$ indicates the range where interactions are mostly homochiral.
@@ E.4 Numerical integration without cutoff
$F_{k_{f}}=\frac{1}{1+\left(\frac{1}{S}\frac{4\pi p_{z}\sqrt{k_{f}^{2}-p_{x}^{2}-p_{z}^{2}}}{L_{y}k_{f}^{3}}\right)^{2}}.$ (140)
@@
Note, finally, that we have neglected 2D-2D interactions in this computation. With isotropic forcing, these should lead to an energy input $T_{\rm 2D-2D}\simeq l_{f}/(4L_{z})\ll 1$ (see [26]), which is negligible for sufficiently large values of $Ro_{\epsilon}$, but not too large. With this hypothesis, the energy injection rate $\epsilon$ corresponds to the energy input in 3D modes only. At low rotation, considering 2D-2D interactions goes beyond our QL theory, as they can disappear due to instabilities, which are not included in our framework.
@@ F.1 Choice of cutoff
The eddy turnover time $\tau_{nl}^{\rm eddy}$ is determined from the Kolmgorov scaling, while the wave-wave interaction time scale $\tau_{nl}^{\rm wave}$ is predicted from wave-turbulent theory [23].
@@
$k_{U}^{\rm wave}$ $=(\Omega p_{z})^{1/3}{{U^{\prime}}}^{2/3}\epsilon^{-1/3},~~~\text{the crossover from condensate- to wave-dominated interactions}$ (157)
$k_{U}^{\rm eddy}$ $={{U^{\prime}}}^{3/2}\epsilon^{-3/2},~~~\text{the crossover from condensate- to eddy-dominated interactions.}$ (158)
These various wavenumbers are explained in the following.
@@
Nonlinear interactions between fluctuation modes occur over a nonlinear time scale $\tau_{nl}$, whose estimation depends on whether mode $\boldsymbol{p}$ is wave-dominated ($p<k_{\Omega}$) or eddy-dominated ($p>k_{\Omega}$). When modes are wave-dominated, they interact with other waves over a slow nonlinear timescale $\tau_{nl}^{\rm wave}$ in (155), which is $\gg\tau_{\Omega}$. In contrast, when modes are eddy-dominated, they interact with each other over the Kolmogorov time scale $\tau_{nl}^{\rm eddy}$ in (154).
@@ F.2 Numerical integration of the Reynolds stress with cutoff
When integrating numerically the Reynolds stress, we consider the Lorentzian filter for the homochiral-wave resonances (140). It yields a smooth transition between the homochiral regime and the mixed-interaction regime around $Ro_{\epsilon}\simeq 0.1$ (see Fig. 9). This near-resonant treatment is more accurate at intermediate $Ro_{\epsilon}$, as it captures the contribution of small near-resonances. However, it fails in the limit $Ro_{\epsilon}\to 0$, as explained in [26].
@@
$\omega_{\boldsymbol{p}}^{+}+\omega_{\boldsymbol{q}}^{-}<\tilde{\omega}{U^{\prime}}$ (162)
in place of (37). This is equivalent to changing $S$ to $\tilde{\omega}S$ in (113). Parameter $\tilde{\omega}$, which must be $\sim 1$, provides an adaptable broadening of the resonant condition. It corresponds to approximating the oscillating factor $\Delta(t)^{s,-s}$ (33) with a parametrized Heaviside function $\Theta(1-|\omega_{\boldsymbol{p}\boldsymbol{q}}^{+-}|/(\tilde{\omega}{U^{\prime}})$. In the results established so far, and in the solution shown in Fig. 4 of the main text, $\tilde{\omega}$ is set to 1. In general, $\tilde{\omega}$ can be used as a tunable parameter of the numerical integration which can better fit the DNS data points. In Figs. 9(a,b), we use $\tilde{\omega}=1.2$. Solutions with $\tilde{\omega}=1$ and $\tilde{\omega}=1.2$ are compared in Fig. 9(c) and (d), respectively.
@@
Figure 10: Energy transfer to the 2D flow, $T_{\rm 3D-2D}/\epsilon$, as a function of $Ro_{\epsilon}$. Theoretical lines show the numerical QL solution with $\tilde{\omega}=1.2$. The black dashed line shows the analytical solution without cutoff, Eq. (152).
The QL computation involving cutoff $K$ is compared to the DNS data points in Figs. 9(c-d). The agreement is quantitatively excellent up to $Ro_{\epsilon}\simeq 0.5$. At large $Ro_{\epsilon}$, the behavior is sensitive to the cutoff and ${U^{\prime}}/\Omega$ decays due to a wave shortage, and reaches zero at $Ro\simeq 0.5$. Our QL result qualitatively reproduces the decay of ${U^{\prime}}/\Omega$ at fixed $Re$, and the transition to strong turbulence seen in the DNS. In addition, we obtain a close quantitative agreement at $Re=23$ and $46$ when using parameter $\tilde{\omega}=1.2$. (Note that we obtain also a good quantitative agreement with the non-cutoff case $\tilde{\omega}=1$ below $Ro_{\epsilon}=1$).